# The occurrence of small, short-period planets younger than 200 Myr with TESS

### 05 Mar 2024 v1

@@ Abstract
Within the first few hundreds of millions of years, many physical processes sculpt the eventual properties of young planets. NASA’s TESS mission has surveyed young stellar associations across the entire sky for transiting planets providing glimpses into the various stages of planetary evolution. Using our own detection pipeline, we search a magnitude-limited sample of 7219 young stars ( Myr) observed in the first four years of TESS for small (), short period ( days) transiting planets. The completeness of our survey is characterized by a series of injection and recovery simulations. Our analysis of TESS 2-minute cadence and Full Frame Image (FFI) light curves recover all known TOIs, as well as four new planet candidates not previously identified as TOIs. We derive an occurrence rate of for mini-Neptunes and for super-Neptunes from the 2-minute cadence data, and % for mini-Neptunes and % for super-Neptunes from FFI data. To independently validate our results, we compare our survey yield with the predicted planet yield assuming Kepler planet statistics. We consistently find a mild increase in the occurrence of super-Neptunes and a significant increase in the occurrence of Neptune-sized planets with orbital periods of days when compared to their mature counterparts. The young planet distribution from our study is most consistent with evolution models describing the early contraction of hydrogen-dominated atmospheres undergoing atmospheric escape and inconsistent with heavier atmosphere models offering only mild radial contraction early on.
@@ 1 Introduction
During the first 100 Myr post-formation, many evolutionary processes sculpt the eventual properties and architectures of young planetary systems. Planet-disc (Lin & Papaloizou, 1986; Ida & Lin, 2004) ( 10 Myr), planet-planet and planet-planetesimal ( 100 Myr) interactions (Chatterjee et al., 2008; Nagasawa et al., 2008; Nagasawa & Ida, 2011; Schlichting et al., 2015) can cause perturbations to planetary orbits, shaping the eventual system stability and architecture. Further, the evolution of the host star and its interactions with evolving planets ( Gyr) can shape their physical properties (e.g., Lammer et al., 2003; Jackson et al., 2012; Owen & Wu, 2013; Ginzburg et al., 2018).
Population studies of mature planetary systems have unveiled scarcities and over-densities within the period-radius distribution of close-in small planets, motivating investigations into the mechanisms sculpting this distribution (e.g. Beaugé & Nesvorný, 2013; Mazeh et al., 2016; Fulton et al., 2017; Hardegree-Ullman et al., 2020; Cloutier & Menou, 2020). Thermal contraction, paired with atmospheric erosion and mass loss driven by photoevaporation (Lopez & Fortney, 2013; Owen & Wu, 2013) and core-powered mass-loss (Ginzburg et al., 2018; Gupta & Schlichting, 2019) are thought to be the dominant processes. Atmospheric erosion driven by photoevaporation resulting from high energy XUV flux is thought to be most active within the first Myr, whereas core-powered mass loss is thought to be a more subtle process predicted to occur over the span of Gyr. However, recent studies have illustrated that the exact timescales and interplay between these two mechanisms are not yet well understood and require further study (Owen & Schlichting, 2023). As a result of these processes, the demographics of the close-in, small-planet population most heavily impacted by these evolutionary processes should vary as a function of age. Kepler (Borucki et al., 2010) and K2 (Howell et al., 2014) identified a number of young planets (e.g., David et al., 2016; Mann et al., 2016, 2018; Rizzuto et al., 2018; Vanderburg et al., 2018; David et al., 2019; Bouma et al., 2022; Barber et al., 2022). These surveys provided the first glimpse into the temporal evolution that small planets undergo early in their lives.
@@
TESS has already detected and characterized tens of young planetary systems (e.g., Newton et al., 2019; Battley et al., 2020; Plavchan et al., 2020; Zhou et al., 2022; Wood et al., 2023), providing initial insights into the differences between young planetary systems and their mature counterparts. Initial investigations into the population of young planets observed by TESS span a wider age range (Fernandes et al., 2023), targeting clusters which already have confirmed planets (Fernandes et al., 2022), and explored methods of light curve extraction that best preserves transit signals in the scenario of heightened photometric variability from stellar activity (Nardiello et al., 2020).
@@ 2 TESS Observations
TESS is an all-sky photometric transit planet survey. TESS observations are conducted via four individual cameras covering a area of the sky, with each sector of observations lasting an average of 27 days. TESS samples the light curves of preselected target stars at 2-minute and 20-second cadences via target pixel stamp observations. These data products are processed by the Science Processing Operation Center (SPOC; Jenkins et al., 2016) at the NASA Ames Research Center. We make use of the SPOC Simple Aperture Photometry (SAP) for our SPOC planet search.
Additionally, TESS observes its entire field of view via Full-Frame Images (FFIs). FFIs were taken every 30-minutes in the TESS primary mission and then every 10-minutes in the first extended mission. The FFIs are downlinked and processed by the MIT Quick-Look Pipeline (QLP; Huang et al., 2020a, b; Kunimoto et al., 2022). We utilize the QLP default aperture raw flux light curves for our QLP planet search program.
We make use of the light curves made available through the Mikulski Archive for Space Telescopes (MAST)111https://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html as of February 2023. This includes the first four years (sector 1-58) of TESS.222We do not make use of newer 200 second FFI light curves in this analysis as they were made available post-February 2023.
@@ 3 Stellar Population
ESA’s Gaia (Gaia Collaboration et al., 2016, 2018, 2021) mission has provided relative- and proper-motions and parallaxes for almost all nearby stars, which has revealed previously unknown comoving populations.
@@
We collate a sample of stars from literature analyses of Gaia kinematics to assemble a population of young stars for our planet search. We make use of associations and clusters assembled from Gagné et al. (2018), Kounkel & Covey (2019), Ujjwal et al. (2020), and Moranta et al. (2022), as well as identified members for well-characterized associations and clusters from Rebull et al. (2016), Rebull et al. (2018), Rebull et al. (2022), and Meingast et al. (2019). Gagné et al. (2018) and Ujjwal et al. (2020) provided updated memberships for previously known young stellar associations. Ujjwal et al. (2020) also updated the ages of the associations via Gaia color-magnitudes and MIST isochrone modeling (Dotter, 2016). Kounkel & Covey (2019) applied unsupervised learning to Gaia DR2 space motions to assemble groups of associated stars, named as Theia groups, and aged via isochrone fits to their color-magnitudes. Moranta et al. (2022) searched for nearby stellar streams within pc using kinematics from Gaia EDR3, named Crius groups. For newly identified groups, Moranta et al. (2022) estimates age ranges for identified associations via isochron fits to the Gaia color-magnitude diagrams. However, we only adopt the members of Crius groups that are affiliated with previously identified associations and clusters that have well-characterized ages. Rebull et al. (2016) and Rebull et al. (2018) use literature photometry and Gaia astrometry to classify previously identified members of the Pleiades cluster and Upper Sco, we make use of stars identified to be best/gold members (see Rebull et al. 2016 Table 1 and Rebull et al. 2018 Table 1). Rebull et al. (2022) selects members of Upper Centarus-Lupus (UCL) and Lower Centaurus-Crux (LCC) from the literature, additionally requiring V and K photometry and no source confusion in TESS; we only adopt members identified as gold members (see Rebull et al. 2022 Table 2). Meingast et al. (2019) used 6D Gaia kinematics to identify overdensities in velocity space, indicative of nearby stellar streams. We make use of the members identified to be part of Pisces-Eridanus stellar stream (Meingast et al., 2019; Curtis et al., 2019). The assigned associations for each source star are listed in Table 1, where a star is listed from more than one source, all references and literature association assignments are noted. Figure 3 shows the galactic coordinates of our stellar population.
@@
We concatenate these membership lists and cross-match them with TICv8 (Stassun et al., 2019). TICv8 stellar parameter values (Stassun et al., 2019) are adopted for all subsequent analyses. With the exception of planet-hosting stars, we do not propagate the stellar parameter uncertainties for individual stars forward in the occurrence rate analyses. We then select for stars with age estimates Myr, stellar radii , stellar masses , effective temperatures K, and TESS band magnitudes . This yields 9847 unique stars, with 8804 having been observed by TESS.
@@
Following previous planet occurrence studies with the Kepler sample (e.g., Berger et al., 2020; Kunimoto & Matthews, 2020), we filter for binaries by removing stars with a GaiaDR3 renormalised unit weight error (RUWE) (Kervella et al., 2022). However, a blind RUWE cut removes stars from our sample that have a resolved companion. Therefore, for any star with RUWE ¿ 1.4, we check for a resolved companion within 2 arcseconds and a delta magnitude of 4, as the resolved companion explains the high RUWE and does not suggest a possible false-positive (El-Badry et al., 2021). If this criterion is matched, the high RUWE star is included in our parent population. This yields 7219 stars with TESS observations. Of these, 7154 were observed in the TESS FFIs and 1927 were preselected to have SPOC 2-minute cadence data. Figure 1 shows the color-magnitude diagram of our parent stellar population.
@@
The population of stars within our SPOC sample is biased toward the proposed science cases for their observations. As a result, the distributions of stellar properties in our SPOC and QLP parent samples differ. Figure 2 shows our QLP sample is skewed towards brighter, earlier-type stars leading to intrinsic differences in our two surveys’ parent populations.
| TIC | R.A. | Dec. | [K] | Parallax [mas] | TESS | Age [Myr] | Association | [days] | ||
@@
Our target list is a collation of multiple catalogs, some of which encompass the traditional clusters and associations, while others are composed of new associations identified via clustering studies. A fraction of our target list may be contamination stars that are not part of their assigned associations, while others may have ages that are older than our thresholds. For example, Kounkel & Covey (2019) estimates a field star contamination rate of 5-10% for the Theia group memberships. From the catalogs we used to collate our parent stellar sample, only Kounkel & Covey (2019) estimates an age uncertainty, dex, which we adopt for our entire parent sample in our forward modeling exercises (see Section 6.1).
@@
To independently verify the contamination rate of our target sample, we measure the rotation periods and variability signatures of our sample with the available TESS photometry. We attempt to derive rotation periods for each star with TESS input catalog effective temperatures K via a Lomb-Scargle (Lomb, 1976; Scargle, 1982) periodogram over all available TESS long cadence photometry from the QLP. We only attempt to derive rotation periods for later type stars in our sample, as they are more likely to host convective envelops and exhibit strong stellar rotational variability at ages up to 200 Myr. The membership contamination rates estimated from the rotational variability of late-type stars are extrapolated and adopted for our entire survey sample (see Section 6.2). We search for stellar rotation periods between 0.5 and 12 days, as we do not attempt to search for rotational periods longer than that of the TESS orbital period. We note that stars with ages Myr should have rotation periods days (e.g. Messina, 2021), and as such the upper limit cutoff should not bias our sample. The rotation periods are then manually vetted to check if the signals resemble that of 1) rotational variability, 2) binarity, 3) non-rotational variability, or 4) no variability. Example light curve figures that were used for our manual vetting process are shown in Appendix D. We analysed rotational signatures for a sample of 6402 stars that satisfy our stellar properties thresholds, of which 4878 exhibit variability resembling rotational modulation from our manual vetting. Our measured rotation periods are presented in Table 1.
@@
Rotation-based age estimates are unreliable for the youngest stars we sample, as pre-main sequence stars that are still contracting exhibit a significant spread in rotation periods for a given age. As such, we do not directly make use of gyrochronological relationships to directly compare against literature-reported isochrone-derived ages. We do make use of the age rotation relationships to quantify the membership contamination rate for our sample.
@@
We make use of the Bouma et al. (2023) gyrochronology relationship to convert rotation periods to age posteriors on a star-to-star basis. This is only performed for the subsample of stars that exhibit rotational variability and are not identified to be binaries based on their light curves and Gaia stellar parameter flags. We find that 9% of stars have gyrochronology-based posteriors with age upper limits Myr, making them inconsistent with the bounds of our sample selection. In addition, late-type young stars should exhibit significant stellar activity due to spot-induced modulations in their light curves. Our manual vetting flagged 30% of G and K stars as not exhibiting significant photometric variations. As such, we conservatively adopt a membership contamination rate ranging over 10-30% for our parent sample in this study.
@@
In addition, we also specifically check for consistency between the literature-reported ages and our gyrochronology-derived ages for the associations that host new planet candidates identified in this study. Figure 4 shows our measured rotation period distributions of LCC, Theia 116, and Theia 214 against that of previously well-characterized populations. These associations host new planet candidates. We find general consistency between their literature ages to the rotation distributions of their members, and that no planet candidates need to be excluded from our occurrence rate due to their parent association being older than 200 Myr.
@@ 4.1 Detrending Stellar Activity
Young stars present additional challenges during the detrending processes of their light curves as a result of heightened activity. Photometric variations on the scale of 1-2% due to rapid stellar rotation, along with stellar flares, require special attention.
@@
4.2 Box Least Squares Transit Detection

### 11 Apr 2024 v2

@@ Abstract
Within the first few hundreds of millions of years, many physical processes sculpt the eventual properties of young planets. NASA’s Transiting Exoplanet Survey Satellite (TESS) mission has surveyed young stellar associations across the entire sky for transiting planets, providing glimpses into the various stages of planetary evolution. Using our own detection pipeline, we search a magnitude-limited sample of 7219 young stars ( Myr) observed in the first 4 yr of TESS for small (), short period ( days) transiting planets. The completeness of our survey is characterized by a series of injection and recovery simulations. Our analysis of TESS 2 minute cadence and Full Frame Image (FFI) light curves recover all known TESS Objects of Interest (TOIs), as well as four new planet candidates not previously identified as TOIs. We derive an occurrence rate of for mini-Neptunes and for super-Neptunes from the 2 minute cadence data, and % for mini-Neptunes and % for super-Neptunes from the FFI data. To independently validate our results, we compare our survey yield with the predicted planet yield assuming Kepler planet statistics. We consistently find a mild increase in the occurrence of super-Neptunes and a significant increase in the occurrence of Neptune-sized planets with orbital periods of days when compared to their mature counterparts. The young planet distribution from our study is most consistent with evolution models describing the early contraction of hydrogen-dominated atmospheres undergoing atmospheric escape and inconsistent with heavier atmosphere models offering only mild radial contraction early on.
@@ 1 Introduction
During the first 100 Myr postformation, many evolutionary processes sculpt the eventual properties and architectures of young planetary systems. Planet-disk (Lin & Papaloizou, 1986; Ida & Lin, 2004, 10 Myr) and planet-planet and planet-planetesimal ( 100 Myr) interactions (Chatterjee et al., 2008; Nagasawa et al., 2008; Nagasawa & Ida, 2011; Schlichting et al., 2015) can cause perturbations to planetary orbits, shaping the eventual system stability and architecture. Further, the evolution of the host star and its interactions with evolving planets ( Gyr) can shape their physical properties (e.g., Lammer et al., 2003; Jackson et al., 2012; Owen & Wu, 2013; Ginzburg et al., 2018).
Population studies of mature planetary systems have unveiled scarcities and overdensities within the period-radius distribution of close-in small planets, motivating investigations into the mechanisms sculpting this distribution (e.g. Beaugé & Nesvorný, 2013; Mazeh et al., 2016; Fulton et al., 2017; Cloutier & Menou, 2020; Hardegree-Ullman et al., 2020). Thermal contraction, paired with atmospheric erosion and mass loss driven by photoevaporation (Lopez & Fortney, 2013; Owen & Wu, 2013) and core-powered mass loss (Ginzburg et al., 2018; Gupta & Schlichting, 2019) are thought to be the dominant processes. Atmospheric erosion driven by photoevaporation resulting from high energy X-ray and UV (XUV) flux is thought to be most active within the first Myr, whereas core-powered mass loss is thought to be a more subtle process predicted to occur over a span of Gyr. However, recent studies have illustrated that the exact timescales and interplay between these two mechanisms are not yet well understood and require further study (Owen & Schlichting, 2023). As a result of these processes, the demographics of the close-in, small planet population most heavily impacted by these evolutionary processes should vary as a function of age. Kepler (Borucki et al., 2010) and K2 (Howell et al., 2014) identified a number of young planets (e.g., David et al., 2016; Mann et al., 2016, 2018; Rizzuto et al., 2018; Vanderburg et al., 2018; David et al., 2019; Barber et al., 2022; Bouma et al., 2022). These surveys provided the first glimpse into the temporal evolution that small planets undergo early in their lives.
@@
TESS has already detected and characterized tens of young planetary systems (e.g., Newton et al., 2019; Battley et al., 2020; Plavchan et al., 2020; Zhou et al., 2022; Wood et al., 2023), providing initial insights into the differences between young planetary systems and their mature counterparts. Initial investigations into the population of young planets observed by TESS span a wider age range (Fernandes et al., 2023), targeting clusters which already have confirmed planets (Fernandes et al., 2022), and explored methods of light-curve extraction that best preserve transit signals in the scenario of heightened photometric variability from stellar activity (Nardiello et al., 2020).
@@ 2 TESS Observations
TESS is an all-sky photometric transit planet survey. TESS observations are conducted via four individual cameras covering a area of the sky, with each sector of observations lasting an average of 27 days. TESS samples the light curves of preselected target stars at 2 minute and 20 s cadences via target pixel stamp observations. These data products are processed by the Science Processing Operation Center (SPOC; Jenkins et al., 2016) at the NASA Ames Research Center. We make use of the SPOC Simple Aperture Photometry (SAP) for our SPOC planet search.
Additionally, TESS observes its entire field of view via Full-Frame Images (FFIs). FFIs were taken every 30 minutes in the TESS primary mission and then every 10 minutes in the first extended mission. The FFIs are downlinked and processed by the MIT Quick-Look Pipeline (QLP; Huang et al., 2020a, b; Kunimoto et al., 2022). We utilize the QLP default aperture raw flux light curves for our QLP planet search program.
We make use of the light curves made available through the Mikulski Archive for Space Telescopes (MAST)111https://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html as of 2023 February. These include the first 4 yr (Sectors 1-58) of TESS.222We do not make use of newer 200 s FFI light curves in this analysis as they were made available post-February 2023.
@@ 3 Stellar Population
ESA’s Gaia (Gaia Collaboration et al., 2016, 2018, 2021) mission has provided relative motions, proper motions, and parallaxes for almost all nearby stars, which has revealed previously unknown comoving populations.
@@
We collate a sample of stars from literature analyses of Gaia kinematics to assemble a population of young stars for our planet search. We make use of associations and clusters assembled from Gagné et al. (2018), Kounkel & Covey (2019), Ujjwal et al. (2020), and Moranta et al. (2022), as well as identified members for well-characterized associations and clusters from Rebull et al. (2016, 2018, 2022), and Meingast et al. (2019). Gagné et al. (2018) and Ujjwal et al. (2020) provided updated memberships for previously known young stellar associations. Ujjwal et al. (2020) also updated the ages of the associations via Gaia magnitudes and MIST isochrone modeling (Dotter, 2016). Kounkel & Covey (2019) applied unsupervised learning to Gaia DR2 space motions to assemble groups of associated stars, named as Theia groups, and aged via isochrone fits to their Gaia magnitudes. Moranta et al. (2022) searched for nearby stellar streams within pc using kinematics from Gaia EDR3, named Crius groups. For newly identified groups, Moranta et al. (2022) estimate age ranges for identified associations via isochrone fits to the Gaia magnitudes. However, we only adopt the members of Crius groups that are affiliated with previously identified associations and clusters that have well-characterized ages. Rebull et al. (2016) and Rebull et al. (2018) use literature photometry and Gaia astrometry to classify previously identified members of the Pleiades cluster and Upper Sco; we make use of stars identified as best/gold members (see Rebull et al. 2016 Table 1; Rebull et al. 2018 Table 1). Rebull et al. (2022) select members of Upper Centarus-Lupus (UCL) and Lower Centaurus-Crux (LCC) from the literature, additionally requiring and photometry and no source confusion in TESS; we only adopt members identified as gold members (see Rebull et al. 2022 Table 2). Meingast et al. (2019) used 6D Gaia kinematics to identify overdensities in velocity space, indicative of nearby stellar streams. We make use of the members identified to be part of Pisces-Eridanus stellar stream (Curtis et al., 2019; Meingast et al., 2019). The assigned associations for each source star are listed in Table 1; where a star is listed from more than one source, all references and literature association assignments are noted. Figure 1 shows the Galactic coordinates of our stellar population.
@@
We concatenate these membership lists and crossmatch them with TICv8 (Stassun et al., 2019). TICv8 stellar parameter values (Stassun et al., 2019) are adopted for all subsequent analyses. With the exception of planet-hosting stars, we do not propagate the stellar parameter uncertainties for individual stars forward in the occurrence rate analyses. We then select for stars with age estimates Myr, stellar radii , stellar masses , effective temperatures K, and TESS band magnitudes (see Figure 2). This yields 9847 unique stars, with 8804 having been observed by TESS.
@@
Following previous planet occurrence studies with the Kepler sample (e.g., Berger et al., 2020; Kunimoto & Matthews, 2020), we filter for binaries by removing stars with a Gaia DR3 renormalized unit weight error (RUWE) (Kervella et al., 2022). However, a blind RUWE cut removes stars from our sample that have a resolved companion. Therefore, for any star with RUWE 1.4, we check for a resolved companion within 2” and a delta magnitude of 4, as the resolved companion explains the high RUWE and does not suggest a possible false positive (El-Badry et al., 2021). If this criterion is matched, the high-RUWE star is included in our parent population. This yields 7219 stars with TESS observations. Of these, 7154 were observed in the TESS FFIs and 1927 were preselected to have SPOC 2 minute cadence data. Figure 3 shows a CMD of our parent stellar population.
@@
The population of stars within our SPOC sample is biased toward the proposed science cases for their observations. As a result, the distributions of stellar properties in our SPOC and QLP parent samples differ. Figure 2 shows our QLP sample is skewed toward brighter, earlier-type stars leading to intrinsic differences in our two surveys’ parent populations.
| TIC | R.A. | Dec. | (K) | Parallax (mas) | TESS | Age (Myr) | Association | (days) | ||
@@
Our target list is a collation of multiple catalogs, some of which encompass the traditional clusters and associations, while others are composed of new associations identified via clustering studies. A fraction of our target list may be contamination stars that are not part of their assigned associations, while others may have ages that are older than our thresholds. For example, Kounkel & Covey (2019) estimate a field star contamination rate of 5-10% for the Theia group memberships. From the catalogs we used to collate our parent stellar sample, only Kounkel & Covey (2019) estimate an age uncertainty, dex, which we adopt for our entire parent sample in our forward modeling exercises (see Section 6.1).
@@
To independently verify the contamination rate of our target sample, we measure the rotation periods and variability signatures of our sample with the available TESS photometry. We attempt to derive rotation periods for each star with TESS input catalog effective temperatures K via a Lomb-Scargle (Lomb, 1976; Scargle, 1982) periodogram over all available TESS long-cadence photometry from the QLP. We only attempt to derive rotation periods for later-type stars in our sample, as they are more likely to host convective envelops and exhibit strong stellar rotational variability at ages up to 200 Myr. The membership contamination rates estimated from the rotational variability of late-type stars are extrapolated and adopted for our entire survey sample (see Section 6.2). We search for stellar rotation periods between 0.5 and 12 days, as we do not attempt to search for rotational periods longer than that of the TESS orbital period. We note that stars with ages Myr should have rotation periods days (e.g. Messina, 2021), and as such the upper limit cutoff should not bias our sample. The rotation periods are then manually vetted to check if the signals resemble that of (1) rotational variability, (2) binarity, (3) nonrotational variability, or (4) no variability. Example light-curve figures that were used for our manual vetting process are shown in Appendix D. We analyzed rotational signatures for a sample of 6402 stars that satisfy our stellar properties thresholds, of which 4878 exhibit variability resembling rotational modulation from our manual vetting. Our measured rotation periods are presented in Table 1.
@@
Rotation-based age estimates are unreliable for the youngest stars we sample, as pre-main-sequence stars that are still contracting exhibit a significant spread in rotation periods for a given age. As such, we do not directly make use of gyrochronological relationships to directly compare against literature-reported isochrone-derived ages. We do make use of the age-rotation relationships to quantify the membership contamination rate for our sample.
@@
We make use of the Bouma et al. (2023) gyrochronology relationship to convert rotation periods to age posteriors on a star-to-star basis. This is only performed for a subsample of stars that exhibit rotational variability and are not identified to be binaries based on their light curves and Gaia stellar parameter flags. We find that 9% of stars have gyrochronology-based posteriors with age upper limits Myr, making them inconsistent with the bounds of our sample selection. In addition, late-type young stars should exhibit significant stellar activity due to spot-induced modulations in their light curves. Our manual vetting flagged 30% of G and K stars as not exhibiting significant photometric variations. As such, we conservatively adopt a membership contamination rate ranging over 10-30% for our parent sample in this study.
@@
In addition, we also specifically check for consistency between the literature-reported ages and our gyrochronology-derived ages for the associations that host new planet candidates identified in this study. Figure 4 shows our measured rotation period distributions of LCC, Theia 116, and Theia 214 against those of previously well-characterized populations. These associations host new planet candidates. We find general consistency between their literature ages to the rotation distributions of their members, and that no planet candidates need to be excluded from our occurrence rate due to their parent association being older than 200 Myr.
@@ 4.1 Detrending Stellar Activity
Young stars present additional challenges during the detrending processes of their light curves as a result of heightened activity. Photometric variations on the scale of 1%-2% due to rapid stellar rotation, along with stellar flares, require special attention.
@@
4.2 Box Least-squares Transit Detection